3 of 29 > Using the graph, calculate the area under the curve as defined by the various points on the curve. 10 Area of A DCE- units2 Area of △ DBF- uni Area of A DAG- units2 about us careers privacy policy terms of use contact us help
Answer #1

Area of a triangle is 1/2 × base × height

Area of triangle DCE is 0.5 × 2.5 × (10-5) = 6.25 units​​2

Area of triangle DBF is 0.5 × 4 × (10-2) = 16 units​​2

Area of triangle DAG is 0.5 × 5 × 10 = 25 units​​2